# CRAFT-GUI: Curriculum-Reinforced Agent For GUI Tasks

### v1

@@ Abstract
As autonomous agents become adept at understanding and interacting with graphical user interface (GUI) environments, a new era of automated task execution is emerging. Recent studies have demonstrated that Reinforcement Learning (RL) can effectively enhance agents’ performance in dynamic interactive GUI environments. However, these methods face two key limitations: (1) they overlook the significant variation in difficulty across different GUI tasks by treating the entire training data as a uniform set, which hampers the agent’s ability to adapt its learning process; and (2) most approaches collapses task-specific nuances into a single, coarse reward, leaving the agent with a uniform signal that yields inefficient policy updates. To address these limitations, we propose CRAFT-GUI, a curriculum learning framework based on Group Relative Policy Optimization (GRPO) that explicitly accounts for the varying difficulty across trajectories. To enable more fine-grained policy optimization, we design a reward function that combines simple rule-based signals with model-judged evaluation, providing richer and more nuanced feedback during training. Experimental results demonstrate that our method achieves significant improvements over previous state-of-the-art approaches, outperforming them by 5.6% on public benchmarks Android Control and 10.3% on our internal online benchmarks, respectively. These findings empirically validate the effectiveness of integrating reinforcement learning with curriculum learning in GUI interaction tasks.
Introduction
@@
Graphical User Interface (GUI) agents offer a promising approach for automating interactions within software applications by autonomously perceiving and manipulating visual elements such as buttons, text fields, and images. Unlike traditional automation tools(Rawles et al. 2023; Wang et al. 2025) that depend on structured inputs or manual scripts, modern GUI agents(Qin et al. 2025; Gou et al. 2024; Lin et al. 2025) leverage recent advances in multimodal large language models to directly interpret GUI screenshots and perform user-specified tasks. This enables robust automation even in environments lacking API or code-level access and paves the way for more accessible and adaptive digital assistants.
@@
Although recent studies have made progress, current GUI agents  (Cheng et al. 2024; Gou et al. 2024; Lu et al. 2025b; Luo et al. 2025; Nong et al. 2024) typically treat all training instances uniformly. However, real-world GUI tasks exhibit substantial variation in difficulty. For instance, as shown in Figure 1, there is a clear distinction between complex tasks and simple tasks. Ignoring these differences in task difficulty leads to several drawbacks: (i) optimization instability during training and (ii) constrained model capability growth. While difficult tasks are included in the training set, the uniform treatment of all samples impedes the agent’s ability to adaptively pivot its learning approach based on task complexity. In other words, learning that disregards knowledge difficulty progression violates established principles of learning – analogous to expecting a child to simultaneously master primary school and university curricula.
@@
Furthermore, when applying reinforcement learning with verified rewards, the reward functions are often based on coarse, rule-based heuristics that fail to differentiate between tasks of varying difficulty. Rather than fundamentally addressing the challenge of how task difficulty impacts learning, current methods  (Lu et al. 2025a) resort to intricate task selection mechanisms and elaborate experience replay strategies. This approach not only adds unnecessary complexity to the training process, but it also neglects to provide the nuanced, task-aware feedback required for efficient policy optimization.
Figure 1: The contrast between complex tasks and simple tasks.
Observing that human users naturally progress from simple to more complex interactions with a GUI, we draw inspiration from this behavior to guide our approach to training GUI agents. Rather than directly transplanting off-the-shelf techniques, our method gradually introduces challenges that align with the intrinsic progression of task difficulty. Previous works in language and vision-language domains—for instance, Kimi k1.5 (Team et al. 2025) and LogicRL (Xie et al. 2025)—have demonstrated that staging learning tasks from basic skills to nuanced reasoning can be beneficial. Similarly, approaches such as MiMo-VL (Xiaomi 2025) refine downstream performance through fine-grained, mixed on-policy reinforcement learning to enhance reward feedback. However, when it comes to GUI automation, defining a clear progression of difficulty is more complicated. The heterogeneity of GUI tasks, with their varying semantic complexity, grounding requirements, and interface dependencies, poses unique challenges that render simple curriculum staging or direct reward modeling insufficient. Given these challenges in GUI environments, we ask an scientific question:
Can a difficulty-aware reinforcement learning strategy boost the performance of GUI tasks?
To tackle these challenges, we propose a curriculum-aware reinforcement learning framework tailored to the unique properties of GUI agent training. Recognizing the inherent heterogeneity and ambiguity of task difficulty in GUI domains, our method organizes training samples based on an adaptive measure that reflects the progression from shorter, simpler trajectories to longer, more complex ones. This sequential learning process enables the agent to gradually acquire more robust and generalizable skills while respecting the nuanced structure of GUI tasks. In addition, we incorporate a fine-grained reward mechanism to provide informative intermediate feedback for key aspects of agent performance, such as tool usage, spatial precision, and semantic correctness. We further integrate understanding-oriented tasks to encourage the agent to jointly develop both low-level action competence and higher-level task comprehension within the reinforcement learning framework. Finally, our experimental results demonstrate that our framework achieves significant improvements over prior state-of-the-art approaches, outperforming them by 5.6% on public benchmarks Android Control and 10.3% on our internal online benchmarks. These findings substantiate the advantages of curriculum-driven reinforcement learning in complex GUI tasks.
In summary, CRAFT-GUI includes the following core contributions:
  1. Curriculum RL Strategy for UI-Agent Model. To address the uneven distribution of task difficulty in GUI scenarios, we propose a phased reinforcement learning strategy that systematically progresses from simpler to more complex tasks based on trajectory characteristics.
  2. Fine-grained Hybrid Reward Mechanisms. We propose a fine-grained reward mechanism that integrates rule-based verification with model-predicted evaluation, enabling progressive and stable convergence of the learning process.
  3. Comprehensive experimental results. We evaluate CRAFT-GUI on both public and private GUI datasets, demonstrating significant improvements over SFT and rule-based RL baselines, with strong performance on both operation and understanding tasks.
Related Work
GUI Agent
@@
Current GUI-Agent implementations for device interaction mainly follow two paths. One relies on closed-source multimodal models(Anthropic 2024; OpenAI 2024) with complex prompt engineering, using visual input and system APIs for perception and decision-making. Representative implementations include the AppAgent (Zhang et al. 2025; Li et al. 2024b; Jiang et al. 2025), series and Mobile-Agent series (Wang et al. 2024b, a).However, these approaches face key limitations: reliance on manual prompt tuning, performance bottlenecks from general-purpose models, and limited adaptability due to the lack of task-specific training for apps, scenarios, or edge cases.
@@
The second approach trains open-source or proprietary models specifically for UI-Agent tasks, emphasizing specialized models for interacting with visual interfaces. Notable implementations from the past year include CogAgent (Hong et al. 2024), AutoGLM (Liu et al. 2024), the Ferret UI series (You et al. 2024; Li et al. 2024c), and UI-TARs (Qin et al. 2025), all adhering to this paradigm. Nevertheless, these methods remain constrained by their dependence on large-scale data for pretraining and fine-tuning, and by performance degradation in out-of-distribution real-world scenarios.
GUI Agent with Reinforcement Learning
@@
The reinforcement learning (RL) training paradigm with verifiable rewards, where reasoning models first generate thinking tokens before producing final result tokens, has emerged as a novel methodology since the introduction of OpenAI o1 (Jaech et al. 2024). Subsequent implementations like DeepSeek-R1 (Guo et al. 2025) achieved remarkable performance through rule-based GRPO training (Shao et al. 2024), providing substantial validation for the test-time scaling paradigm. In domains involving data processing, code generation, and logical reasoning (particularly GUI operation), where verifiable reward signals can be readily derived through predefined rules, rule-based RL training strategies are gaining increasing adoption within the GUI-Agent modeling community. UI-R1 (Lu et al. 2025b) and GUI-R1 (Luo et al. 2025) applying rule-based GRPO training for operation-oriented tasks; InfiGUI-R1 (Liu et al. 2025a) implementing supervised fine-tuning (SFT) cold-start prior to GRPO training; UIShift (Gao, Zhang, and Xu 2025) restricting GRPO training exclusively to the LLM component; ARPO (Lu et al. 2025a) proposing an experience replay-optimized RL algorithm.
@@
However, current GUI-Agent implementations exhibit three fundamental limitations: 1) Data Distribution Bias: None address the detrimental impact of imbalanced difficulty distributions in training data through systematic data decoupling approaches. 2) Task Scope Restriction: Excessive focus on operation-oriented tasks with inadequate support for understanding-oriented tasks, hindering the development of more generally intelligent GUI-Agent systems. 3) Reward Signal Granularity: Overly simplistic reward formulations lacking multi-dimensional, fine-grained reward decomposition constrain further performance improvements.
Method
@@
In this section, we present the key components of our method. We first revisit Group Relative Policy Optimization (GRPO) (Shao et al. 2024) and describe how it is combined with curriculum learning to improve training stability and sample efficiency in complex GUI environments. We then introduce our task-specific reward design, followed by a detailed formulation of the training procedure,
Curriculum-Guided GRPO Training
@@
We adopt Group Relative Policy Optimization (GRPO) as the foundational reinforcement learning algorithm in our framework. Unlike traditional methods such as PPO(Schulman et al. 2017), GRPO eliminates the need for a separate value function by leveraging group-wise advantage estimation, substantially reducing memory usage and computational overhead. This makes it particularly suitable for GUI-based training under limited hardware resources.
@@
Input: initial policy model $\pi_{\theta_{\text{init}}}$; reward function $r_{\varphi}$; curriculum datasets $\mathcal{C}$
Output: optimized policy $\pi_{\theta}$
Initialize policy model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$;
for $\mathcal{D}\in\mathcal{C}=\{\mathcal{D}_{\text{basic}},\mathcal{D}_{\text{intermediate}},\mathcal{D}_{\text{advanced}},\mathcal{D}_{\text{ui}}\}$ do
   Set reference model $\pi_{\text{ref}}\leftarrow\pi_{\theta}$;
 for step =1=1 to $M$ do
      Sample batch $\mathcal{B}\sim\mathcal{D}$;
      Update old policy: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$;
    foreach question $q\in\mathcal{B}$ do
         Sample $G$ outputs $\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$;
         Compute rewards $\{r_{i}\}_{i=1}^{G}=r_{\varphi}(o_{i},q)$;
         Compute advantages $\hat{A}_{i}$ using group relative estimation;
      end foreach
     Update $\pi_{\theta}$ by maximizing GRPO objective;
   end for
end for
@@
Fine-Grained Verified Reward Design
To enable a more generalizable and capable UI agent, we design task-specific reward functions for both mobile operation and visual understanding tasks. Inspired by the rule-based reward paradigm introduced in DeepSeek-R1(Guo et al. 2025)—known for its simplicity and effectiveness—we adopt verified reward mechanisms that combine format validation and outcome assessment. This approach has recently expanded from LLMs to VLMs, as demonstrated in works such as VLM-R1 (Shen et al. 2025), and Visual-RFT (Liu et al. 2025b). In our framework, most rewards are defined through verifiable rules, providing lightweight yet reliable signals to guide curriculum-based reinforcement learning.
Mobile Operation Tasks
@@
Tool Type Reward. In mobile operation tasks, we follow Qwen2.5VL (Bai et al. 2025)’s agent paradigm by formulating action spaces through agent tool call: click tool, swipe tool, type tool, long press tool, open app tool, system button tool, wait tool, and terminate tool. This toolkit achieves near-comprehensive coverage of terminal operation requirements. If the predicted tool $o_{tool}$ matches the ground truth tool ${gt}_{tool}$, it is assigned with a reward of 1, or 0 otherwise. As shown in Equation 3:
@@
Overlong Response Penalty During GRPO training we observe a progressive increase in response length, primarily driven by expanding thinking token sequences. If left unregulated, this phenomenon risks triggering explosive thinking token growth that may ultimately lead to model performance collapse. Inspired by DAPO (Yu et al. 2025)’s optimization principles, we implement an adaptive length constraint mechanism. This approach applies graduated penalties to responses exceeding predetermined length thresholds, effectively suppressing overgeneration tendencies while preserving essential reasoning components. As shown in Equation 6, $L_{\max}$ is the max output length, and $L_{cache}$ is the buffer length for soft penalty.
@@
Visual Understanding Tasks
@@
Experiments
Implement Details
@@
Evaluation benchmarks and metrics. We established an online evaluation benchmark to comprehensively assess model capabilities in real-world operational environments, focusing on three critical dimensions: page comprehension accuracy, contextual reasoning effectiveness, and integrated decision-making proficiency. To validate the methodological efficacy proposed in this work, parallel evaluations were conducted on open-source benchmark suites including Android Control(Li et al. 2024a) and AndroidWorld(Rawles et al. 2024). The primary evaluation metric employed is task success rate (SR), calculated as Equation 8:
@@
Online Benchmark for Operation Task
Figure 3: Domain-specific Distribution of Our Online Benchmark.
Experimental Results
@@
Table 1: Training Data Overview
Complexity Training Data Details
Task Type Samples (k) Device
Basic Level Operation 21 iPhone Android iPad
Intermediate Level Operation 28
Advanced Level Operation 8
Understanding 23
Model Mobile Application Categories Average
Food-delivery Dine-in Medical Finance Insurance Gaming
Claude-3.7-Sonnet 31.7% 28.3% 44.4% 24.2% 25.6% 61.7% 35.9%
GPT-4.1 19.5% 18.8% 32.4% 21.3% 18.0% 56.2% 27.7%
UI-TARS-72B 65.4% 57.2% 67.5% 61.8% 55.1% 85.3% 65.4%
Qwen2.5-VL-72B 65.7% 52.3% 67.6% 57.4% 56.9% 88.6% 64.8%
CRAFT-GUI-7B 43.9% 48.4% 74.5% 67.4% 54.2% 96.4% 64.1%
CRAFT-GUI-32B 76.5% 73.8% 73.5% 77.5% 60.4% 92.7% 75.7%
In-house Datasets Training and Evaluation
We developed a proprietary training dataset comprising six major categories of mobile applications: food delivery, in-store dining, medical services, financial services, insurance services, and gaming apps. This multi-domain corpus supports phased curriculum learning through progressive complexity escalation. Details can be found in Table 1.
Table 2 presents the comparative evaluation results between two CRAFT-GUI variants (varying in model scale) and the top-tier industry model such as Claude-3.7-Sonnet(Anthropic 2024) and GPT-4.1(OpenAI 2025) on our online mobile operation benchmark and visual understanding benchmark. The histograms demonstrate CRAFT-GUI’s statistically significant superiority across all six mobile application categories following multi-phase reinforcement learning optimization. Additional curriculum learning results for understanding tasks are in the appendix.

### v2

@@ Abstract
As autonomous agents become adept at understanding and interacting with graphical user interface (GUI) environments, a new era of automated task execution is emerging. Recent studies have demonstrated that Reinforcement Learning (RL) can effectively enhance agents’ performance in dynamic interactive GUI environments.
@@
However, these methods face two key limitations: (1) they overlook the significant variation in difficulty across different GUI tasks by treating the entire training data as a uniform set, which hampers the agent’s ability to adapt its learning process; and (2) most approaches collapse task-specific nuances into a single, coarse reward, leaving the agent with a uniform signal that yields inefficient policy updates. To address these limitations, we propose CRAFT-GUI, a curriculum learning framework based on Group Relative Policy Optimization (GRPO) that explicitly accounts for the varying difficulty across trajectories. To enable more fine-grained policy optimization, we design a reward function that combines simple rule-based signals with model-judged evaluation, providing richer and more nuanced feedback during training. Experimental results demonstrate that our method achieves significant improvements over previous state-of-the-art approaches, outperforming them by 7.1% on public benchmarks AndroidWorld and 10.3% on our private constructed dataset, respectively. These findings empirically validate the effectiveness of integrating reinforcement learning with curriculum learning in GUI interaction tasks.
1 Introduction
@@
Graphical User Interface (GUI) agents offer a promising approach for automating software interactions by autonomously perceiving and manipulating visual elements. Unlike traditional automation tools Rawles et al. (2023); Wang et al. (2025) that depend on structured inputs, modern GUI agents Qin et al. (2025); Gou et al. (2024); Lin et al. (2025) leverage multimodal large language models to directly interpret GUI screenshots and perform user-specified tasks, enabling robust automation even without API access.
@@
Although recent studies have made progress, current GUI agents Cheng et al. (2024); Gou et al. (2024); Lu et al. (2025b); Luo et al. (2025); Nong et al. (2024) typically treat all training instances uniformly. However, real-world GUI tasks exhibit substantial variation in difficulty, as shown in Figure 1. Ignoring these differences leads to two critical drawbacks: (i) optimization instability during training and (ii) constrained model capability growth. Uniform treatment of samples impedes the agent’s ability to adaptively adjust its learning based on task complexity—analogous to expecting a child to simultaneously master elementary and university curricula.
Figure 1: Contrast between complex and simple GUI tasks.
@@
Furthermore, when applying reinforcement learning with verified rewards, existing reward functions rely on coarse, rule-based heuristics that fail to differentiate task difficulty. Rather than fundamentally addressing how task difficulty impacts learning, current methods Lu et al. (2025a) resort to intricate task selection and experience replay strategies, adding unnecessary complexity while neglecting nuanced, task-aware feedback for efficient policy optimization.
Inspired by how humans naturally progress from simple to complex GUI interactions, we propose a curriculum-aware reinforcement learning framework tailored to GUI agent training. While previous works Team et al. (2025); Xie et al. (2025); Xiaomi (2025) have shown benefits of staged learning in language domains, GUI automation presents unique challenges: heterogeneity in semantic complexity, grounding requirements, and interface dependencies render simple curriculum staging insufficient. This motivates our research question: Can a difficulty-aware reinforcement learning strategy boost GUI task performance?
Our method organizes training samples based on trajectory complexity, progressing from shorter, simpler tasks to longer, more complex ones. We incorporate fine-grained reward mechanisms providing informative feedback for tool usage, spatial precision, and semantic correctness. By integrating both operation and understanding tasks, the agent jointly develops low-level action competence and high-level task comprehension. Experimental results demonstrate significant improvements: 7.1% gain on AndroidWorld and 10.3% on our private benchmark over prior state-of-the-art approaches.
Contributions. (1) A curriculum RL strategy that systematically progresses from simple to complex tasks based on trajectory characteristics. (2) Fine-grained hybrid reward mechanisms integrating rule-based and model-predicted evaluation for stable convergence. (3) Comprehensive evaluation demonstrating strong improvements over SFT and RL baselines on both operation and understanding tasks.
2 Related Work
2.1 GUI Agent
@@
Current GUI agent implementations follow two main paradigms. The first relies on closed-source multimodal modelsAnthropic (2024); OpenAI (2024) with prompt engineering, using visual input and system APIs for perception and decision-making. Representative works include AppAgent Zhang et al. (2025); Li et al. (2024b); Jiang et al. (2025) and Mobile-Agent series Wang et al. (2024b; a). However, these approaches face key limitations: reliance on manual prompt tuning, performance bottlenecks from general-purpose models, and limited adaptability to diverse scenarios.
@@
The second approach trains specialized models for GUI tasks, emphasizing task-specific optimization. Notable implementations include CogAgent Hong et al. (2024), AutoGLM Liu et al. (2024), Ferret-UI series You et al. (2024); Li et al. (2024c), UI-Venus  Gu et al. (2025) and UI-TARS Qin et al. (2025). While these methods demonstrate improved performance on specific benchmarks, they remain constrained by dependence on large-scale pretraining data and performance degradation in out-of-distribution scenarios.
2.2 Reinforcement Learning for GUI Agents
@@
The RL paradigm with verifiable rewards has emerged as a promising approach, inspired by reasoning models like OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025). In GUI domains where verifiable reward signals can be derived through predefined rules, RL-based training strategies are gaining increasing adoption. UI-R1 Lu et al. (2025b), GUI-G2  Tang et al. (2025) and GUI-R1 Luo et al. (2025) apply rule-based GRPO training for operation tasks; InfiGUI-R1 Liu et al. (2025a) implements SFT cold-start prior to GRPO; UIShift Gao et al. (2025) restricts GRPO training to the LLM component; ARPO Lu et al. (2025a) proposes experience replay-optimized algorithms.
@@
However, current implementations exhibit three fundamental limitations: (1) Data distribution bias: none address the detrimental impact of imbalanced difficulty distributions through systematic curriculum approaches; (2) Task scope restriction: excessive focus on operation tasks with inadequate support for understanding tasks; (3) Reward granularity: overly simplistic reward formulations lacking multi-dimensional feedback. Our work addresses these gaps through curriculum-based RL with fine-grained, task-aware reward mechanisms.
3 Method
@@
In this section, we present the key components of our method. We first revisit Group Relative Policy Optimization (GRPO) Shao et al. (2024) and describe how it is combined with curriculum learning to improve training stability and sample efficiency in complex GUI environments. We then introduce our task-specific reward design, followed by a detailed formulation of the training procedure.
3.1 Curriculum-Guided GRPO Training
@@
We adopt Group Relative Policy Optimization (GRPO) as the foundational reinforcement learning algorithm in our framework. Unlike traditional methods such as PPOSchulman et al. (2017), GRPO eliminates the need for a separate value function by leveraging group-wise advantage estimation, substantially reducing memory usage and computational overhead. This makes it particularly suitable for GUI-based training under limited hardware resources.
@@
Input: initial policy model $\pi_{\theta_{\text{init}}}$; reward functions $r_{\varphi}$; task prompts $\mathcal{D}$
Output: $\pi_{\theta}$
1 policy model $\pi_{\theta}\leftarrow\pi_{\theta_{\text{init}}}$;
2
3for each dataset $\mathcal{D}$ in curriculum order do
4   reference model $\pi_{\text{ref}}\leftarrow\pi_{\theta}$;
5
6 for $\text{step}=1,\dots,M$ do
7      Sample a batch $\mathcal{D}_{b}$ from $\mathcal{D}$;
8
9     Update the old policy model $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$;
10
11     Sample $G$ outputs $\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{b}$;
12
13     Compute rewards $\{r_{i}\}_{i=1}^{G}$ for each sampled output $o_{i}$ by running $r_{\varphi}$;
14
15     Compute $\hat{A}_{i}$ for $o_{i}$ with ground truth $gt_{i}$ through group relative advantage estimation;
16
17     Update the policy model $\pi_{\theta}$ by maximizing the GRPO objective;
18
19   end for
20
21 end for
@@ 17     Update the policy model $\pi_{\theta}$ by maximizing the GRPO objective;
3.2 Fine-Grained Verified Reward Design
To enable a more generalizable and capable UI agent, we design task-specific reward functions for both mobile operation and visual understanding tasks. Inspired by the rule-based reward paradigm introduced in DeepSeek-R1Guo et al. (2025)—known for its simplicity and effectiveness—we adopt verified reward mechanisms that combine format validation and outcome assessment. This approach has recently expanded from LLMs to VLMs, as demonstrated in works such as VLM-R1 Shen et al. (2025), and Visual-RFT Liu et al. (2025b). In our framework, most rewards are defined through verifiable rules, providing lightweight yet reliable signals to guide curriculum-based reinforcement learning.
3.2.1 Mobile Operation Tasks
@@
Tool Type Reward. In mobile operation tasks, we follow Qwen2.5VL Bai et al. (2025)’s agent paradigm by formulating action spaces through agent tool call: click tool, swipe tool, type tool, long press tool, open app tool, system button tool, wait tool, and terminate tool. This toolkit achieves near-comprehensive coverage of terminal operation requirements. If the predicted tool $o_{tool}$ matches the ground truth tool ${gt}_{tool}$, it is assigned with a reward of 1, or 0 otherwise. As shown in Equation 3:
@@
Overlong Response Penalty During GRPO training we observe a progressive increase in response length, primarily driven by expanding thinking token sequences. If left unregulated, this phenomenon risks triggering explosive thinking token growth that may ultimately lead to model performance collapse. Inspired by DAPO Yu et al. (2025)’s optimization principles, we implement an adaptive length constraint mechanism. This approach applies graduated penalties to responses exceeding predetermined length thresholds, effectively suppressing overgeneration tendencies while preserving essential reasoning components. As shown in Equation 6, $L_{\max}$ is the max output length, and $L_{cache}$ is the buffer length for soft penalty.
@@
3.2.2 Visual Understanding Tasks
@@
4 Experiments
4.1 Implement Details
@@
Evaluation benchmarks and metrics. We established an online evaluation benchmark to comprehensively assess model capabilities in real-world operational environments, focusing on three critical dimensions: page comprehension accuracy, contextual reasoning effectiveness, and integrated decision-making proficiency. To validate the methodological efficacy proposed in this work, parallel evaluations were conducted on open-source benchmark suites including Android ControlLi et al. (2024a) and AndroidWorldRawles et al. (2024). The primary evaluation metric employed is task success rate (SR), calculated as Equation 8:
@@
4.2 Experimental Results
@@
4.2.1 Open-source Datasets Training and Evaluation
4.2.2 In-house Datasets Training and Evaluation
We developed a proprietary training dataset comprising 80K samples across six major categories of mobile applications: food delivery, in-store dining, medical services, financial services, insurance services, and gaming apps. All data are collected from multiple device platforms (iPhone, Android, iPad) to ensure cross-platform generalization. This multi-domain corpus is organized into three difficulty levels to support phased curriculum learning: Basic level (21K operation samples, $\leq$3 steps), Intermediate level (28K operation samples, 4-8 steps), and Advanced level (8K operation samples and 23K understanding samples, >>8 steps or involving visual reasoning tasks).
Our method significantly outperforms industry baselines including Claude-3.7-Sonnet Anthropic (2024) and GPT-4.1 OpenAI (2025) across all application domains. CRAFT-GUI-32B achieves 75.7% average success rate, representing 10.3% improvement over the best baseline (detailed results in Appendix Table 4). Progressive curriculum learning yields consistent gains: our 32B model improves from 62.6% (Stage 1) to 75.7% (Stage 3), while the 7B variant progresses from 43.9% to 64.1% (see Appendix Table 5 for stage-wise breakdown). Additional understanding task results are provided in Appendix Table 7.
4.3 Ablation Study